OpenStudy (anonymous):

The function f(x) = 4(3)x represents the growth of a dragonfly population every year in a remote swamp. Erin wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for Erin's purpose, and what is the new growth rate?

OpenStudy (solomonzelman):

you probably meant \(\normalsize\color{royalblue}{ \rm f(x)=4(3)^x }\)

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

where f(0) is the initial population. f(1) is the population after year 1. f(2) is the population after year 2. Now, Erin wants a formula that will be calculating the growth every quarter of the year.

OpenStudy (solomonzelman):

I will give you an example of what I would do if I wanted to manipulate our formula so that it will be giving the growth every half of a year. \(\normalsize\color{royalblue}{ \rm f(x)=4(3)^{x/2} }\) and now, f(0) => initial population f(1) => population after 1/2 of a year f(2) => population after 2/2 of a year (or 1 year) f(3) => population after 3/2 of a year (or 1 year and a half) f(4) => population after 4/2 of a year (or 2 years) f(5) => population after 5/2 of a year (or 2 years and a half) f(4) => population after 6/2 of a year (or 3 years)

OpenStudy (anonymous):

whaat? im so confused :c

OpenStudy (solomonzelman):

do you have something specific to ask, or you just don't get anything from what I posted (not that I am claiming to be anything like a very good tutor)

OpenStudy (anonymous):

how does half a year have to do with her wanting to know 4 times a year?

OpenStudy (solomonzelman):

I gave you example how to do a formula that tells you a growth after 1/2 of a year.

OpenStudy (solomonzelman):

just an example, not answer....

OpenStudy (anonymous):

these are my options f(x) = 4(3)x; growth rate 300% f(x) = 4(3)x, growth rate 4% f(x) = 4(1.32)x; growth rate 4% f(x) = 4(1.32)4x; growth rate 32%

OpenStudy (solomonzelman):

ok

OpenStudy (solomonzelman):

I would like to see your attempt to generate the requested formula please/

OpenStudy (anonymous):

so she wants 4*(1.32)^x?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm f(x)=4(3)^x }\) \(\normalsize\color{royalblue}{ \rm f(x)=4(3)^{x/4} }\) \(\normalsize\color{royalblue}{ \rm f(x)=4(\sqrt[4]{3})^{x} }\) \(\normalsize\color{black}{ \rm \sqrt[4]{3}\approx 1.32 }\) \(\normalsize\color{royalblue}{ \rm f(x)=4(1.32)^{x} }\) (answer)

OpenStudy (anonymous):

so the growth rate is 32 %?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

because you are taking 132% of the entire previous amount when multiplying times 1.32. That means you are adding another 32% to the 100% you had already.

OpenStudy (anonymous):

so f(x) 3(1.32)^4x with a growth rate of 32?

OpenStudy (solomonzelman):

f(x)=... correct

OpenStudy (anonymous):

would you mind helping me with another one ??

OpenStudy (solomonzelman):

I am with another student right now, but when I finish and if I have more time, I will attempt my best to help you.